OpenStudy (anonymous):

Carmen is saving money to buy a game. so far she has saved $36,which is three-fourths of the total cost of the game. how much does the game cost

OpenStudy (whpalmer4):

You can write this as the following equation: \[\$36 = \frac{3}{4} X\]where \(X\) is the cost of the game. Do you know how to solve that equation?

OpenStudy (anonymous):

i dont

OpenStudy (shaik0124):

x *(3/4)=36 x*3=36*4 now divide both sides by 3 u will get the game cost

OpenStudy (whpalmer4):

Ok, I'll walk you through it. First thing to know about solving equations is that if you do the same thing to both sides, the equation remains true. For example, if you and your friend both have $10 (each), you have equal amounts of money. If someone gives each of you another $5, you now have a different amount of money than when you started, but the equality is still there — you both have the same amount of money as the other person. Make sense?

OpenStudy (whpalmer4):

You want to manipulate that equation so that you end up with just the variable, \(X\), on one side of the \(=\) sign, and everything else on the other side.

OpenStudy (whpalmer4):

As you have a fraction there, a good first step is to multiply both sides of the equation by the denominator of the fraction (the number on the bottom). This will get rid of the fraction: \[\$36 = \frac{3}{4}X\]\[4*\$36 = 4*\frac{3}{4}X\]\[$144 = \cancel{4}\frac{3}{\cancel{4}}X\]\[\$144 = 3X\] What might you do next to get X alone?